What is the molarity of a sodium hydroxide solution if 9.0 mL of the solution is titrated to the end point with 10.0 mL of standard 0.20M hydrochloric acid

1
Expert's answer
2022-03-10T10:44:56-0500

Moles of HCL

0.20mol were in 1000mL of solution

How many moles could be in 10.0mL of the solution

"=\dfrac{0,2mol x 10.0mL}{1000mL} =0.002mol"


Equation for the reaction

NaOH + HCl "\to" NaCl + H2O


Mole ratio of HCl:NaOH is 1:1


Moles of NaOH use is therefore;

"=\dfrac{1}{1}x0.002mol = 0.002mol"


Concentration of NaOH "= \dfrac{0.002molx1000ml}{9.0mL} = 0.222M"